OpenStudy (anonymous):

calculate the amount of heat gnerated while transferring 90000 coloumbs of charge between the two terminals of a battery of 40v in one hour. also find the power expended in the process

OpenStudy (anonymous):

@paki

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

\[H = V \times Q = V \times I \times t\]

OpenStudy (anonymous):

like i didvided 90000 by 40 and i got 2250 is dat rightt

Parth (parthkohli):

I think you should multiply those two instead of dividing.

OpenStudy (anonymous):

oh noooo

OpenStudy (anonymous):

u mean 36000000 i sthe answer

Parth (parthkohli):

Yeah, just not that many zeroes.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

how du find powerr

Parth (parthkohli):

If you want to find power, then divide the heat generated by time for which the heat was generated.

OpenStudy (anonymous):

thnxxx